# Robot Pouring: Identifying Causes of Spillage and Selecting Alternative Action Parameters Using Probabilistic Actual Causation

### v1

@@ 4.2.3 Example 3
Figure 8(c) shows the reference probability and the probabilities obtained for contrastive $RD$ values obtained from inequality (9). In the area where probability raising holds, the contrastive probabilities show a sharp decrease around $RD\approx 1.0$ and a low probability for $RD>1.1$. We select the alternative values $RD=0.99$ (chance-level probability) and $RD=1.02$ (low probability $\approx 0.2$). The results from 100 replications are shown in Figure 8(f) and Figure 8(h). The replications with the alternative parameters show that $RD=1.02$ is better suited to avoid spillage.
Figure 8(d) shows the reference probabilities and the probabilities obtained for contrastive FU values obtained from inequalities (10) and (11). In the area where probability raising holds, we observe low probabilities for $FU<0.6$. For $FU>0.6$, we observe smooth probability increments. We select the alternative values $FU=0.75$ (chance-level probability) and $FU=0.64$ (low probability $\approx 0.2$). The results from 100 replications are shown in Figure 8(j) and Figure 8(l). The replications with the alternative parameters show that $FU=0.64$ is better suited to avoid spillage.
@@ 4.2.4 Example 4
Figure 9(d) shows the reference probabilities and the probabilities obtained for contrastive FU values obtained from inequalities (10) and (11). In the area where probability raising holds, we observe low probabilities for $FU<0.4$. For $FU>0.4$, we observe smooth probability increments. We select the alternative values $FU=0.76$ (chance-level probability) and $FU=0.53$ (low probability $\approx 0.2$). The results from 100 replications are shown in Figure 9(f) and Figure 9(h). The replications with the alternative parameters show that $FU=0.53$ is better suited to avoid spillage.
@@ 6 Conclusions
The reliability of the analysis of actual causation relies on the correctness of the causal graph structure and the estimated conditional probability distributions. This constitutes a major challenge for implementing an actual causation analysis in real-life tasks as it requires 1) a careful selection of the variables’ representation, 2) determining the structure of the causal graph, and 3) estimating conditional probability distributions. In this work, we used a realistic simulation of the pouring task to cover an ample combinatorial space of task parameters, which would have been cumbersome to replicate in a real environment. The simulation provided us with a large dataset to learn the causal structure of the task using a causal discovery algorithm and to estimate its conditional probability distributions using neural networks. These methods leverage the availability of a large data sample. Alternatively, the causal structure can be determined using expert knowledge, and the conditional probabilities can be estimated using traditional methods (e.g., regression).
@@ 6.1 Causal discovery using the PC algorithm - Tetrad setup and bootstrapping results
To validate the causal relationships inferred from the data, we performed a bootstrapping analysis. We ran the PC algorithm on 1000 bootstraps, producing 1000 different structures. The edge-type frequencies obtained from bootstrapping indicate whether the discovered causal relationships are stable across different bootstrap samples [Glymour et al., 2019]. In Figure 13, we report the frequency of the edge types between variables. For ease of interpretation and comparison, the frequency of edge type is reported as a proportion of the number of bootstraps. Given the different edge types (including “no edge”), we interpret an edge frequency larger than 0.5 as stable. The reported DAG, termed discovered DAG includes the stable edges and no-edges. The PC algorithm discovered direct edges, edges that are definitely direct (DD), and undirected edges (i.e., the data are consistent with $X\rightarrow Y$ and $X\leftarrow Y$ [Malinsky and Danks, 2017]).

### v3

@@ 5.2.3 Example 3
Figure 8(c) shows the reference probability and the probabilities obtained for contrastive $RD$ values obtained from inequality (9). In the area where probability raising holds, the contrastive probabilities show a sharp decrease around $RD\approx 1.0$ and a low probability for $RD>1.1$. We select the alternative values $RD=0.99$ (chance-level probability) and $RD=1.02$ (low probability $\approx 0.2$), as shown in Figures 8(e) and 8(g), respectively. The results from 100 replications are shown in Figure 8(f) and Figure 8(h). The replications with the alternative parameters show that $RD=1.02$ is better suited to avoid spillage.
Figure 8(d) shows the reference probabilities and the probabilities obtained for contrastive FU values obtained from inequalities (10) and (11). In the area where probability raising holds, we observe low probabilities for $FU<0.6$. For $FU>0.6$, we observe smooth probability increments. We select the alternative values $FU=0.75$ (chance-level probability) and $FU=0.64$ (low probability $\approx 0.2$), as shown in Figures 8(i) and 8(k), respectively. The results from 100 replications are shown in Figure 8(j) and Figure 8(l). The replications with the alternative parameters show that $FU=0.64$ is better suited to avoid spillage.
@@ 5.2.4 Example 4
Figure 9(d) shows the reference probabilities and the probabilities obtained for contrastive FU values obtained from inequalities (10) and (11). In the area where probability raising holds, we observe low probabilities for $FU<0.4$. For $FU>0.4$, we observe smooth probability increments. We select the alternative values $FU=0.76$ (chance-level probability) and $FU=0.53$ (low probability $\approx 0.2$), as shown in Figures 9(e) and 9(g), respectively. The results from 100 replications are shown in Figure 9(f) and Figure 9(h). The replications with the alternative parameters show that $FU=0.53$ is better suited to avoid spillage.
@@
5.3 Evaluation of alternative actions to prevent spillage
In this section, we evaluate the capabilities of the analysis of probabilistic actual causation to guide the selection of alternative parameters to prevent spillage. For this evaluation, we generated a test dataset of 3000 pouring trials. The trial parameters of the test dataset were sampled from the same distributions used for the training dataset, as described in Table 2. Spillage occurred in 1216 trials and 1784 trials were successful.
Following steps described in Section 2.1, we conducted the analysis of probabilistic actual causation on the spillage trials. We identified the range of contrastive values where probability raising holds, and within these values, we selected the subset of values with probability of spillage$<0.1$. From this subset, the closest value to the current parameter was selected as alternative parameter. We ran the trial using the alternative parameter and recorded the outcome.
As explained in Section 2.1, the analysis of actual causation can be applied to different variables, one at a time. When inequality (1) does not hold for any of the contrastive values, the variable cannot be regarded as an actual cause. This indicates that there are no alternative values for the analyzed variable. Even if inequality (1) holds, it can also occur that the probabilities within the range of contrastive values where probability raising holds lay above the chance level or above the desired probability threshold (0.1 in our case). Therefore, it may happen that no alternative values for the variable being analyzed can be identified. To evaluate this aspect, we conducted the analysis of probabilistic actual causation on the $RC$, $FU$, and $RD$ variables, and determined the percentage of the spillage trials for which an alternative parameter satisfying the probability threshold could be identified. For $RC$, an alternative parameter could be identified in 2.7% of the spillage trials, for $FU$ in 59% of the spillage trials, and for $RD$ in 97.9% of the spillage trials.
The marked differences between variables can be attributed to the causal structure of the task and the magnitude of the causal probabilities presented in Section 5.1. $RC$ has an indirect effect on $S$ through $RV$. $FU$ also has an indirect effect on $S$ through $RV$ and a direct effect. Comparing the causal probabilities $P(S|do(RC))$ (Figure 5(a)), with $P(S|do(FU))$ and $P(S|do(FU,RV))$ (Figures 5(b) and 5(c), respectively), it can be observed that the effect of $FU$ on $S$ is stronger than the effect of $RC$. Under these considerations, finding an alternative $RC$ value is less likely than finding an alternative $FU$ value. In contrast to $RC$ and $FU$, $RD$ has a direct effect on $S$, and the $P(S|do(RD))$ (Figure 5(e)) shows a range with low spillage probability, which leads to a higher likelihood of finding an alternative value.
[…]
Finally, we evaluate the empirical success rates observed when conducting trial replications using the alternative $FU$ and $RD$ values. For this, we selected a random subset of 100 spillage trials from the test dataset. We ran 10 replications of each trial of the subset using the alternative $RD$ or $FU$ values. As a result, we obtained an empirical success rate for each trial ($\frac{number\ of\ successful\ replications}{10}$). The histogram of the empirical success rates obtained with the alternative $RD$ values is shown in Figure 10(a), and the histogram obtained for the alternative $FU$ values is shown in Figure 10(b). In general, the observed success rates are in line with the probability criterion used to select the alternative values. Nevertheless, lower success rates of 0.6 and 0.7 were also obtained for some trials. Crucially, all the success rates are above the chance level, which provides empirical support for the usefulness of the alternative values.
(a) Empirical success rates using alternative $RD$ values.
(b) Empirical success rates using alternative $FU$ values.
Figure 10: Empirical success rates obtained over 10 replications.
@@ 7 Conclusions
The reliability of the analysis of actual causation relies on the correctness of the causal graph structure and the estimated conditional probability distributions. This constitutes a major challenge for implementing an actual causation analysis in real-life tasks as it requires 1) a careful selection of the variables’ representation, 2) determining the structure of the causal graph, and 3) estimating conditional probability distributions. The methods described in Section 4 constitute state-of-the art best common practices to obtain reliable and robust causal modeling results. Specifically, we used a realistic simulation of the pouring task to cover an ample combinatorial space of task parameters, which would have been cumbersome to replicate in a real environment. The simulation provided us with a large dataset to learn the causal structure of the task using a causal discovery algorithm with bootstrapping and to estimate its causal probability distributions using neural networks. In addition to the information provided in Section 4, in the supplementary material we discuss further modeling assumptions and we provide empirical support to the correctness of the causal model.
@@ 7.1 Causal discovery using the PC algorithm - Tetrad setup and bootstrapping results
To validate the causal relationships inferred from the data, we performed a bootstrapping analysis. Conducting a bootstrapping analysis is recommended for causal discovery[Malinsky and Danks, 2017, Glymour et al., 2019]. We ran the PC algorithm on 1000 bootstraps, producing 1000 different structures. The edge-type frequencies obtained from bootstrapping indicate whether the discovered causal relationships are stable across different bootstrap samples [Glymour et al., 2019]. In Figure 14, we report the frequency of the edge types between variables. For ease of interpretation and comparison, the frequency of edge type is reported as a proportion of the number of bootstraps. Given the different edge types (including “no edge”), we interpret an edge frequency larger than 0.5 as stable. The reported DAG, termed discovered DAG includes the stable edges and no-edges. The PC algorithm discovered direct edges, edges that are definitely direct (DD), and undirected edges (i.e., the data are consistent with $X\rightarrow Y$ and $X\leftarrow Y$ [Malinsky and Danks, 2017]).
@@
The reported DAG was constructed from the edges with frequencies larger than 0.5. As shown in Figure 14, the discovered DAG results from edge frequencies larger than 0.75. This increases our confidence in the correctness of the discovered causal structure, which is crucial for the analysis of probabilistic actual causation.
Internally, we also conducted the causal discovery analysis with bootstrapping using Tetrad’s GFCI, FGES, GRASP, and BOSS algorithm implementations (for all algorithms, we used the Degenerate Gaussian Likelihood Ratio Test (DG-LRT)). The edge frequencies slightly vary across algorithms, but the discovered structure is the same. For simplicity, we opted to report the PC results since it is a well-established algorithm with extensively studied properties [Malinsky and Danks, 2017, Glymour et al., 2019].
7.2 Assumption of no latent variables
The PC algorithm assumes no latent confounders or unobserved variables. In this respect, we rely on our analysis of the task and the data-generating process to support this assumption.
Based on the analysis of the task described in Section “4.2 DAG Variables" we assume that the variables used to represent the data-generating process of the simulation capture all the relevant causes of spillage. The variables represent the stochastic effect of 1) the containers’ characteristics (capacities and diameters, expressed as the variables $RC$ and $RD$) and 2) the poured amount (fullness and volume, expressed as the variables $FU$ and $RV$) on the probability of spilling (variable $S$). The randomness of the outcome $S$ results from the interplay between $RC$, $RD$, $FU$, and $RV$ and the behavior of the particles during the pouring movement. The parameters the physics engine uses (particle size and density) to simulate the particles’ behavior have a constant value. Therefore, they are not included as DAG variables.
Additionally, the pouring movement was executed with constant rotation velocity and angle. Thus, these were not included as DAG variables. Based on these considerations, we are confident that we included all the relevant random variables of the data-generating process of the simulation and assume that there are no latent confounders or other unobserved variables with a causal effect on spillage.
@@ 7.3 Implementation of the neural autoregressive density estimators
The hyper-parameters were selected based on the results reported by Garrido et al. [2021]. In their extensive analysis of the performance of different hyper-parameters (hidden layers, number of units, and learning rates), they conclude that no single combination of hyper-parameters is superior to others. We selected two hidden layers with 16 units. We noticed that increasing the number of layers and units did not reduce the loss or improve the prediction performance of the network. Based on this empirical observation, which is in line with the observations of Garrido et al. [2021], we abstained from conducting any further systematic search or comparison of the hyper-parameter space.
7.4 Empirical support to the correctness of the causal model
The analysis of probabilistic actual causation relies on the correctness of the causal structure and the estimated causal probabilities. It is important to note that the causal structure of the data-generating process is unknown. Therefore, no ground truth causal graph is available to benchmark the structure and the causal probabilities learned from the training dataset.
In simulation, we have control over the parameters used in each pouring trial. The actual movement of the particles during the pouring movement, which determines whether or not spillage occurs, depends on the interaction between the trial parameters and the physics simulation. This interaction between the trial parameters and the spillage outcome is unknown. Therefore, the benchmarking of the causal graph can only be conducted on the level of evaluating the spillage predictions against the ground truth. We evaluate the spillage predictions on a test dataset of 3000 pouring trials. The trial parameters of the test dataset were sampled from the same distribution used for the training dataset (see description in Section “4.2 DAG Variables"). For each trial, we compute the causal probability of spillage $P(S|do(FU,RC,RD,RV))$. If $P(S|do(FU,RC,RD,RV))\geq 0.5$, the outcome prediction is labeled as $S=True$. We compare the predicted outcome with the actual outcome. The prediction results are summarized in the confusion matrix shown in Figure 15. Among the spillage trials, the causal model predictions yield 90.8% true positives and 9.2% false negatives. On the other hand, the causal model predictions yield 94.5% true negatives and 5.5% false positives among the no-spillage trials. These results provide empirical support that the causal model corresponds with the ground truth data-generating process.
Figure 15: Confusion matrix.